OpenStudy (anonymous):

The length of a rectangle is 3 times the width. THe perimeter is 96 cm. Find the width and length

OpenStudy (anonymous):

\[ \text{Say we have as follows...}\\\ \ \ w - \text{the width of the rectangle}\\\ \ \ \ l\ - \text{the length of the rectangle}\\\text{We're given as follows...}\\\ \ \ l=3w\\\ \ \ 2(l+w)=96\\\text{Now, we need to eliminate a variable to solve for the other.}\\\ \ \ 2(3w+w)=96\\\ \ \ 4w=48\\\ \ \ w=12\\\text{Since we have solved for the width, we can now determine the length.}\\\ \ \ l=3(12)=36 \]